There is an electric field at point P, A very small charge isplaced at this point and experiences a force. Another very small charge is then placed at this point and experiences a force that differs in both magnitude and direction from that experienced by the first charge. How can these two different forces result from the single electric charge that exsists at point P?


1
Expert's answer
2021-07-19T15:40:53-0400

Gives

Point P electric field E then charge q

"E=\frac{kq}{r^2}"

Another point contain charge q1 at point P

"F_1=\frac{kqq_1}{r^2}=q_1E\rightarrow(1)"

Again a different poin charge q2 exist at point P

"F_2=\frac{kqq_2}{r^2}=q_2E\rightarrow(2)"

equation (1) and (2) are different directions and magnitude at same point P